OpenStudy (anonymous):

A video game club charges a fixed annual membership fee of $19 and $5 per video game rented. Let f(n) represent the total annual cost of renting n video games. Which of the following functions best represents the relationship between f(n) and n if the membership was decreased by $10 the next year?

OpenStudy (anonymous):

f(n) = 19n + 5 f(n) = 5n + 9 f(n) = 4n + 29 f(n) = 5n + 29

hartnn (hartnn):

what do u think ?

hartnn (hartnn):

$5 per video game, and n video games, so total (5n) $s so there should be 5n term in your equation

hartnn (hartnn):

membership fees $19, but DECREASED by 10$ this year, so now its ?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

i mean B!!

hartnn (hartnn):

you meant b ?

hartnn (hartnn):

yes! B is correct! :)

OpenStudy (anonymous):

yes lol idk why i put d ;-;

hartnn (hartnn):

haha! you always do that .... first put incorrect option, then say i meant *the correct option* :P

OpenStudy (fatimab2000):

the real answer is d